anzamukosha:

Lea walked1/2mile in 1/3 hour. How long will it take Lea to walk 2 1/4 miles? Enter your answer as a mixed number in simplest form in the box.

xOliver2:

I think I know this.

xOliver2:

\[2~~~1/4=9/4\] \[27/12\div9/12=1~~1/2\] 1 1/2 hours